OpenStudy (anonymous):

(These questions are the ones i got wrong on my test I need someone to help me answer them please.) Solve the system of equations. y=-0.04x^2+8.3x y=1.12x+1.6 B. (0.22,1.85) C. (179.28, 202.39) D. (0.22,1.85) and (179.28, 202.39)

OpenStudy (anonymous):

Note that you have two equations, each of which equals y. So you can set the right side of each equation equal to each other: \[\begin{align} -.04x^2+8.3x=1.12x+1.6\\ .04x^2-7.18x+1.6=0\\ x=.22, \;179.28\\ y=1.12x+1.6\\ y=1.85,\;202.39 \end{align}\] So your answers are (0,22, 1.85) and (179.28, 202.39): D.

OpenStudy (phi):

notice that you if you set the right hand sides equal to each other (see above), or subtract the two equations, you end up with a quadratic equation (has an x^2 term). Generally speaking, you will end up with 2 x values (and so 2 y values also). Looking at the the possible answers (if in a hurry), notice that only the last choice gives two answers for (x,y). So I would go with that one. If I had a calculator, I would try replacing x with 0.22 in the 2nd equation (it looks simpler) to see if you get 1.85. Then test with x= 179.28, to see if we get 202.39 If you have lots of time, use the quadratic formula to solve for the 2 x values, and then find the y values.